OpenStudy (anonymous):

How do I simplify this expression: √3 / √3 + √2? The answer is 3 - √6... how?

OpenStudy (anonymous):

is sqrt(2) in the denominator?

OpenStudy (anonymous):

no, there's no root 2 in the denominator

OpenStudy (ash2326):

\(\sqrt 3/( \sqrt 3+ \sqrt 2)\) now multiply numerator and denominator by \(\sqrt 3- \sqrt 2\) we have \( \sqrt 3/( \sqrt 3 + \sqrt 2) *( \sqrt 3 - \sqrt 2)/( \sqrt 3 - \sqrt 2)\) we get \[ (3-2 \sqrt 6)/( 3-2)\] or \[3-2\sqrt 6\]

OpenStudy (ash2326):

sorry it'd be \[3-\sqrt 6\]

OpenStudy (anonymous):

why is it 3 - 2√6?

OpenStudy (ash2326):

ignore the 2 int the last two steps

OpenStudy (ash2326):

did you get??

OpenStudy (anonymous):

but why is it 3 - √6? Did you cancel something?

OpenStudy (ash2326):

when we multiplied \(\sqrt 3 -\sqrt 2\) in the numerator and denominator , denominator became 1 . and numerator became \[3-\sqrt 6\]

OpenStudy (anonymous):

can you please show me? I'm still confused about what you just said.. so sorry D:

OpenStudy (ash2326):